# MS-TP-25-28

### v2

@@ Abstract
A massive vector field is a highly promising candidate for dark matter in the universe. A salient property of dark matter is its negligible or null coupling to ordinary matter, with the exception of gravitational interaction. This poses a significant challenge in producing the requisite amount of dark particles through processes within the Standard Model. In this study, we examine the production of a vector field during inflation due to its direct interaction with the inflaton field through kinetic and axion-like couplings as well as the field-dependent mass. The gradient-expansion formalism, previously proposed for massless Abelian gauge fields, is extended to include the longitudinal polarization of a massive vector field. We derive a coupled system of equations of motion for a set of bilinear functions of the vector field. This enables us to address the nonlinear dynamics of inflationary vector field production, including backreaction on background evolution. To illustrate this point, we apply our general formalism to a low-mass vector field whose kinetic and mass terms are coupled to the inflaton via the Ratra-type exponential function. The present study investigates the production of its transverse and longitudinal polarization components in a benchmark inflationary model with a quadratic inflaton potential. It has been demonstrated that pure mass coupling is able to enhance only the longitudinal components. By turning on also the kinetic coupling, one can get different scenarios. As the coupling function decreases, the primary contribution to the energy density is derived from the transverse polarizations of the vector field. Conversely, for an increasing coupling function, the longitudinal component becomes increasingly significant and rapidly propels the system into the strong backreaction regime.
@@ I Introduction
According to observational data, dark matter (DM) is the most abundant form of matter in the Universe, accounting for roughly one quarter of its total energy density. The existence of DM is supported by numerous independent astrophysical observations, including galaxy rotation curves, gravitational lensing, the dynamics of hot gas in clusters, matter clustering during large-scale structure formation, primordial nucleosynthesis, and measurements of the cosmic microwave background (CMB). All of these indicate the presence of an invisible, massive component that interacts primarily through gravity [1, 2, 3, 4]. Data from primordial nucleosynthesis and the cosmic microwave background further establish that DM must be non-baryonic in nature. The Standard Model (SM) of particle physics cannot account for the totality of DM: the only viable SM candidates, neutrinos, are too light (constituting hot DM) and are tightly constrained by structure formation and CMB data to contribute no more than a few percent of the total DM abundance. Thus, the composition of the dominant DM component remains unknown, making its nature a central open problem in both cosmology and particle physics. A wide range of extensions of the SM propose various DM candidates [5, 6, 7, 8, 9, 10, 11, 12, 13, 14]. Among them, the most widely studied are weakly interacting massive particles (WIMPs), axions, and dark photons, as they are also motivated by considerations beyond DM. Other possibilities include primordial black holes, self-interacting DM, fuzzy DM, asymmetric DM, and $Q$-balls (for a recent comprehensive review, see Ref. [3]).
@@
A very light dark photon with a small but nonzero mass can serve as a dark matter candidate if it is generated non-thermally in the early Universe. In one class of scenarios [27, 28], the dark photon mass is generated through the Stückelberg mechanism, which requires a non-minimal coupling to gravity. When the Hubble parameter decreases below the dark photon mass, the field begins to oscillate, and these oscillations behave as non-relativistic matter, effectively acting as cold dark matter. An alternative possibility is that dark photons are produced from vacuum fluctuations during inflation [29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49]. Unlike the electromagnetic field, a massive vector field can be generated without requiring a nonminimal coupling to gravity or the inflaton. This is possible because the presence of a mass term in the Lagrangian automatically breaks conformal invariance of the action — a property that standard electromagnetism does not possess. However, in such cases, only the longitudinal modes are efficiently generated; see, e.g., Refs. [29, 35, 38]. To achieve substantial production of transverse modes, one must introduce an interaction either with the spacetime curvature [50, 34, 37, 43, 47, 46, 45] or with the inflaton field [51, 52, 33, 36, 31, 53, 42, 41, 44], as is commonly done in analogous scenarios of inflationary magnetogenesis (for a review, see Refs. [54, 55, 56]).
In this work, we focus on the latter mechanism and consider a single massive Abelian vector field interacting with the inflaton field via generic kinetic (or dilatonic) coupling, axial coupling, as well through the inflaton-dependent mass. These couplings are included simultaneously in order to cover a large range of scenarios considered in the literature. We are paying special attention to the regime of strong backreaction of inflationary massive vector field pro- duction.
At the same time, we neglect the kinetic mixing of the vector DM field with the ordinary electromagnetic field. In this way, we can study the production of the dark vector field fully decoupled from the SM sector. Inclusion of the kinetic mixing is an important next step which we intend to address in the future work.
@@
The remainder of the paper is organized as follows. In Sec. II, we introduce our model of vector dark matter coupled to the inflaton field during inflation. Section III is devoted to the evolution of the transverse polarization components of the vector field, analyzed both in Fourier space and in position space; in the latter case, we derive the equations of the GEF and supplement them with the appropriate boundary terms. In Sec. IV, we carry out a similar analysis for the longitudinal polarization component of the vector field. All resulting equations of motion are summarized in Sec. V, where we also discuss the couplings between different sectors and their impact on the inflationary background. In Sec. VI, we specify the coupling functions within a simple benchmark scenario and present numerical results for several parameter choices. For the case of negligible backreaction, we validate the GEF against the mode-by-mode solution in Fourier space. We then explore the regime of strong backreaction. Finally, in Sec. VII, we provide our conclusions and outline directions for future research. In Appendix A we discuss in more detail the equations of motion for the longitudinal polarization of the vector field.
@@ III.1 Mode evolution in Fourier space
As time evolves, the other two terms in brackets become more and more important. Finally, at some point, they become comparable to unity. In the physically interesting situation, the whole expression in brackets can cross zero, and the mode becomes tachyonically unstable. We define the momentum of the mode which crosses the horizon in the following way:
@@ III.2 Gradient-expansion formalism
If the produced gauge field is strong enough so that its backreaction becomes non-negligible, the dynamics of the system becomes highly nonlinear and all Fourier modes are coupled. In this case, it is more convenient to keep working in the coordinate space and introduce the following set of bilinear quantities:
@@
This relation gives a unique result in the transverse sector. Using Eq. (45), we can easily get the explicit expressions for the bilinear quantities with negative indices:
@@
Here $\mathcal{H}_{\rm b}$ is the magnetic helicity.
Now, using the Proca equations (13)–(15), it is straightforward to derive equations of motion for the bilinear quantities (42)–(44):
@@
These equations are supplemented by extra terms on the right-hand side, the so-called boundary terms, which originate from the contributions of vacuum gauge-field modes crossing the horizon. They act like a quantum source pumping the energy from vacuum modes and, by this, at least partially counter-acting the redshift due to the universe’s expansion.
@@
Here, the upper integration boundary is replaced by a finite momentum. By this, we regularize the UV divergent integrals for the bilinear quantities. Although this is far from a rigorous renormalization (as it is done in quantum field theory in Minkowski space), this trick allows us to separate the relevant Fourier modes and disregard the contributions of vacuum-like modes.
@@
where $k_{h}=k_{h}(t)$ is the wavenumber of the mode crossing the horizon at the given time $t$, see Eq. (35) for its explicit form. The value of $X$ depends on time for two reasons. First, its spectral density $dX/d\ln k$ changes over time. And secondly, the integral over momentum has a variable upper integration boundary. Terms on the left-hand side in Eqs. (III.2)–(51) take into account the first time dependence. In order to take care of the second one, we need to introduce the boundary terms:
@@
To calculate this boundary term, we need to know the spectral density of the quantity $X$ at the moment the mode crosses the horizon. For this, we use the explicit expressions for the mode functions $\mathcal{A}_{\lambda}$ and $\mathcal{D}_{\lambda}$ given in Eqs. (39) and (III.1) which are valid close to the moment of horizon crossing.
@@
Now, using Eq. 56, we obtain the expressions for the boundary terms:
@@
and similar relations for $\mathcal{G}^{(n_{\mathrm{max}}+1)}$ and $\mathcal{B}^{(n_{\mathrm{max}}+1)}$. Note that more complicated truncation conditions, like the one proposed in Ref. [63], can be used as well.
Despite the fact that Eqs. (III.2) and (III.2) for $n$th-order quantities also contain the terms of the order (n1)(n-1), the system is not infinite in the direction of negative orders. Indeed, Eq. (51) does not contain such lower-order terms, and this allows one to close the system. Only three negative-order quantities are relevant: $\mathcal{G}^{(-1)}$, $\mathcal{B}^{(-1)}$, and $\mathcal{B}^{(-2)}$.
@@ IV.1 Mode evolution in Fourier space
Equation (62) gives the relation between the mode functions $\mathcal{A}_{L}(t,k)$ and $\mathcal{D}_{L}(t,k)$:
@@
which constitutes, together with the previous one, the simplest system of equations suitable for the numerical implementation. In Appendix A, we derive the boundary conditions which supplement the system above; they are based on the assumption that the longitudinal polarization of the gauge field is in the Bunch–Davies vacuum state far inside the Hubble horizon. If the parameter $\mu=m/(\sqrt{I_{1}}H)\ll 1$ (which is the case for the absolute majority of the dark photon models), the boundary conditions take the form:
@@
One can also derive a single second-order differential equation for the mode function $\mathcal{A}_{L}$, in full analogy with the transverse-polarizations case; cf. Eq. (III.1). However, in this case, it takes a rather cumbersome form, which we give in the Appendix A [see Eq. (A) there] not to distract the reader from the main flow of the paper. One can use this equation for numerical calculations as well.
@@
From Eq. (74) we can find a reasonable definition of the momentum of the mode which crosses the horizon. In fact, when the second term in brackets becomes comparable to unity, the behavior of the mode starts deviating from the vacuum solution. Therefore, we define
@@
where again the upper monotonic envelope is used to ensure that once the mode crosses the horizon it never crosses it back during inflation. Further, for any given mode $k$, we can find a moment of time $t_{h,L}$ when it crosses the horizon:
@@ IV.2 Gradient-expansion formalism
Now, using either the definitions of functions $\mathcal{Q}^{(2p)}$, $\mathcal{F}^{(2p)}$, $\mathcal{K}^{(2p+1)}$ in coordinate space, Eqs. (87)–(89), or the spectral representations above, we use Eqs. (85)–(86) or, equivalently, Eqs. (66)–(67) in order to derive the equations of motion for the bilinear functions:
@@
The system of equations of the GEF is also infinite, as in the transverse case. Indeed, equations for $\mathcal{Q}^{(2p)}$ and $\mathcal{F}^{(2p)}$ contain $\mathcal{K}^{(2p+1)}$ while the equation for the latter contains $\mathcal{Q}^{(2p+2)}$ and $\mathcal{F}^{(2p+2)}$ etc. It can be truncated in two possible ways:
@@ V Full system of equations
The quantities constructed from the magnetic field are the simplest ones since they contain only the transverse components:
@@
$=\frac{m^{4}a^{2}}{I_{1}}\langle A_{0}(-\triangle)A_{0}\rangle=\frac{m^{2}}{I_{1}}\mathcal{Q}^{(-2)}\,$, (104)
where in the last line we used Eq. (15) for the divergence of the electric field.
Further, we express the terms quadratic in the vector potential. The square of the 0-component simply reads
@@
The square of the 3-vector decomposes similarly to the electric field contribution:
@@ VI.1 Model specification
Initial conditions for the inflaton field are set at roughly 61 $e$-fold from the end of inflation (without gauge fields) by setting $\phi(0)=\phi_{0}=15.5\,M_{\rm P}$; the corresponding value for the inflaton derivative is taken from the slow-roll attractor solution $\dot{\phi}(0)\approx-V^{\prime}M_{\rm P}/\sqrt{3V}\big|_{\phi_{0}}=-\sqrt{2/3}m_{\phi}M_{\rm P}\approx-4.9\times 10^{-6}\,M_{\rm P}^{2}$.
@@
This allows us to explicitly omit all terms containing the $m^{2}/I_{1}$ factor in the Friedmann equations (V)–(V), Klein–Gordon equation (V) as well as in the GEF systems of equations (III.2)–(51) and (93)–(IV.2). Note that this does not correspond to switching to a massless vector field and discarding its longitudinal modes. We will see that if the mass is small but time-dependent (in fact, inflaton-dependent), there may be efficient vector-field production, especially of its longitudinal modes. In other words, although we are setting $\mu\ll 1$, we keep $M=\dot{m}/(mH)$ to be finite.
In the low-mass approximation above, the transverse GEF system (III.2)–(51) takes the form which is exactly the same as in the case of a massless gauge field. It was extensively studied in the literature previously, see Refs. [57, 58, 60, 65, 63]. Note that effects of time-dependent mass do not enter the set of equations for transverse modes; they are only sensitive to the kinetic and axial couplings. On the other hand, the longitudinal modes of the vector field are totally insensitive to the axial coupling as it never enters the corresponding equations of motion. Moreover, the kinetic coupling only enters the boundary terms while also not appearing in the GEF system (93)–(IV.2) explicitly. Since we are primarily interested in studying the production of longitudinal modes of comparable magnitude to the transverse ones, in our numerical analysis, we keep only the mass couplings to the inflaton and in some cases also the kinetic coupling. Thus, we always set $I_{2}(\phi)=0$ in the remaining part of the paper.
For the kinetic coupling, we assume the functional dependence proposed by Ratra [68] in the context of inflationary magnetogenesis:
@@
where $\beta$ is the dimensionless coupling constant of the model. In principle, in order to avoid the strong coupling problem, one should take only the values $\beta>0$. However, in our numerical analysis below, we will consider also some negative values just to see the qualitative differences in the picture of the vector-field production. [^3] In a realistic setup which avoids the strong coupling problem, one can accommodate the coupling function with negative $\beta$ at finite time intervals if they are followed by the intervals with $\beta>0$ (e.g., in a case of a piecewise coupling function); see, e.g., Refs. [69, 70, 71, 72, 73] for the case of inflationary magnetogenesis.

### v3

@@ Abstract
A massive vector field is a highly promising candidate for dark matter in the Universe. A salient property of dark matter is its negligible or null coupling to ordinary matter, with the exception of gravitational interaction. This poses a significant challenge in producing the requisite amount of dark particles through processes within the Standard Model. In this study, we examine the production of a vector field during inflation due to its direct interaction with the inflaton field through kinetic and axionlike couplings as well as the field-dependent mass. The gradient-expansion formalism, previously proposed for massless Abelian gauge fields, is extended to include the longitudinal polarization of a massive vector field. We derive a coupled system of equations of motion for a set of bilinear functions of the vector field. This enables us to address the nonlinear dynamics of inflationary vector field production, including backreaction on background evolution. To illustrate this point, we apply our general formalism to a low-mass vector field whose kinetic and mass terms are coupled to the inflaton via the Ratra-type exponential function. The present study investigates the production of its transverse and longitudinal polarization components in a benchmark inflationary model with a quadratic inflaton potential. It has been demonstrated that pure mass coupling is able to enhance only the longitudinal components. By turning on also the kinetic coupling, one can get different scenarios. As the coupling function decreases, the primary contribution to the energy density is derived from the transverse polarizations of the vector field. Conversely, for an increasing coupling function, the longitudinal component becomes increasingly significant and rapidly propels the system into the strong backreaction regime.
@@ I Introduction
According to observational data, dark matter (DM) is the most abundant form of matter in the Universe, accounting for roughly one quarter of its total energy density. The existence of DM is supported by numerous independent astrophysical observations, including galaxy rotation curves, gravitational lensing, the dynamics of hot gas in clusters, matter clustering during large-scale structure formation, primordial nucleosynthesis, and measurements of the cosmic microwave background (CMB). All of these indicate the presence of an invisible, massive component that interacts primarily through gravity [1, 2, 3, 4]. Data from primordial nucleosynthesis and the cosmic microwave background further establish that DM must be nonbaryonic in nature. The Standard Model (SM) of particle physics cannot account for the totality of DM: the only viable SM candidates, neutrinos, are too light (constituting hot DM) and are tightly constrained by structure formation and CMB data to contribute no more than a few percent of the total DM abundance. Thus, the composition of the dominant DM component remains unknown, making its nature a central open problem in both cosmology and particle physics. A wide range of extensions of the SM propose various DM candidates [5, 6, 7, 8, 9, 10, 11, 12, 13, 14]. Among them, the most widely studied are weakly interacting massive particles, axions, and dark photons, as they are also motivated by considerations beyond DM. Other possibilities include primordial black holes, self-interacting DM, fuzzy DM, asymmetric DM, and $Q$-balls (for a recent comprehensive review, see Ref. [3]).
@@
A very light dark photon with a small but nonzero mass can serve as a dark matter candidate if it is generated nonthermally in the early Universe. In one class of scenarios [27, 28], the dark photon mass is generated through the Stückelberg mechanism, which requires a nonminimal coupling to gravity. When the Hubble parameter decreases below the dark photon mass, the field begins to oscillate, and these oscillations behave as nonrelativistic matter, effectively acting as cold dark matter. An alternative possibility is that dark photons are produced from vacuum fluctuations during inflation [29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49]. Unlike the electromagnetic field, a massive vector field can be generated without requiring a nonminimal coupling to gravity or the inflaton. This is possible because the presence of a mass term in the Lagrangian automatically breaks conformal invariance of the action — a property that standard electromagnetism does not possess. However, in such cases, only the longitudinal modes are efficiently generated; see, e.g., Refs. [29, 35, 38]. To achieve substantial production of transverse modes, one must introduce an interaction either with the spacetime curvature [50, 34, 37, 43, 47, 46, 45] or with the inflaton field [51, 52, 33, 36, 31, 53, 42, 41, 44], as is commonly done in analogous scenarios of inflationary magnetogenesis (for a review, see Refs. [54, 55, 56]).
In this work, we focus on the latter mechanism and consider a single massive Abelian vector field interacting with the inflaton field via generic kinetic (or dilatonic) coupling, axial coupling, as well through the inflaton-dependent mass. These couplings are included simultaneously in order to cover a large range of scenarios considered in the literature. We are paying special attention to the regime of strong backreaction of inflationary massive vector field production.
At the same time, we neglect the kinetic mixing of the vector DM field with the ordinary electromagnetic field. In this way, we can study the production of the dark vector field fully decoupled from the SM sector. Inclusion of the kinetic mixing is an important next step, which we intend to address in future work.
@@
The remainder of the paper is organized as follows. In Sec. II, we introduce our model of vector dark matter coupled to the inflaton field during inflation. Section III is devoted to the evolution of the transverse polarization components of the vector field, analyzed both in Fourier space and in position space; in the latter case, we derive the equations of the GEF and supplement them with the appropriate boundary terms. In Sec. IV, we carry out a similar analysis for the longitudinal polarization component of the vector field. All resulting equations of motion are summarized in Sec. V, where we also discuss the couplings between different sectors and their impact on the inflationary background. In Sec. VI, we specify the coupling functions within a simple benchmark scenario and present numerical results for several parameter choices. For the case of negligible backreaction, we validate the GEF against the mode-by-mode solution in Fourier space. We then explore the regime of strong backreaction. Finally, in Sec. VII, we provide our conclusions and outline directions for future research. In the Appendix A we discuss in more detail the equations of motion for the longitudinal polarization of the vector field.
@@ III.1 Mode evolution in Fourier space
As time evolves, the other two terms in brackets become more and more important. Finally, at some point, they become comparable to unity. In the physically interesting situation, the whole expression in brackets can cross zero, and the mode becomes tachyonically unstable. We define the momentum of the mode that crosses the horizon in the following way:
@@ III.2 Gradient-expansion formalism
If the produced gauge field is strong enough so that its backreaction becomes nonnegligible, the dynamics of the system becomes highly nonlinear and all Fourier modes are coupled. In this case, it is more convenient to keep working in the coordinate space and introduce the following set of bilinear quantities:
@@
This relation gives a unique result in the transverse sector. Using Eq. (45), we can easily get the explicit expressions for the bilinear quantities with negative indices,
@@
Here $\mathcal{H}_{B}$ is the magnetic helicity.
Now, using the Proca equations (13)–(15), it is straightforward to derive equations of motion for the bilinear quantities (42)–(44),
@@
These equations are supplemented by extra terms on the right-hand side, the so-called boundary terms, which originate from the contributions of vacuum gauge-field modes crossing the horizon. They act like a quantum source pumping the energy from vacuum modes and, by this, at least partially counteract the redshift due to the Universe’s expansion.
@@
Here, the upper integration boundary is replaced by a finite momentum. By this, we regularize the UV divergent integrals for the bilinear quantities. Although this is far from a rigorous renormalization (as it is done in quantum field theory in Minkowski space), this trick allows us to separate the relevant Fourier modes and disregard the contributions of vacuumlike modes.
@@
where $k_{h}=k_{h}(t)$ is the wave number of the mode crossing the horizon at the given time $t$, see Eq. (35) for its explicit form. The value of $X$ depends on time for two reasons. First, its spectral density $dX/d\ln k$ changes over time. And second, the integral over momentum has a variable upper integration boundary. Terms on the left-hand side in Eqs. (III.2)–(51) take into account the first time dependence. In order to take care of the second one, we need to introduce the boundary terms,
@@
To calculate this boundary term, we need to know the spectral density of the quantity $X$ at the moment the mode crosses the horizon. For this, we use the explicit expressions for the mode functions $\mathcal{A}_{\lambda}$ and $\mathcal{D}_{\lambda}$ given in Eqs. (39) and (III.1), which are valid close to the moment of horizon crossing.
@@
Now, using Eq. 56, we obtain the expressions for the boundary terms,
@@
and similar relations for $\mathcal{G}^{(n_{\mathrm{max}}+1)}$ and $\mathcal{B}^{(n_{\mathrm{max}}+1)}$. Note that more complicated truncation conditions, like the one proposed in Ref. [63], can be used as well. In practice, the required truncation order is determined numerically: we increase the order $n_{\mathrm{max}}$ until the solution converges and no longer changes with further increases of $n_{\mathrm{max}}$.
Despite the fact that Eqs. (III.2) and (III.2) for $n$th-order quantities also contain the terms of the order (n1)(n-1), the system is not infinite in the direction of negative orders. Indeed, Eq. (51) does not contain such lower-order terms; this allows one to close the system. Only three negative-order quantities are relevant: $\mathcal{G}^{(-1)}$, $\mathcal{B}^{(-1)}$, and $\mathcal{B}^{(-2)}$.
@@ IV.1 Mode evolution in Fourier space
Equation (62) gives the relation between the mode functions $\mathcal{A}_{L}(t,k)$ and $\mathcal{D}_{L}(t,k)$,
@@
which constitutes, together with the previous one, the simplest system of equations suitable for the numerical implementation. In Appendix A, we derive the boundary conditions which supplement the system above; they are based on the assumption that the longitudinal polarization of the gauge field is in the Bunch–Davies vacuum state far inside the Hubble horizon. If the parameter $\mu=m/(\sqrt{I_{1}}H)\ll 1$ (which is the case for the absolute majority of the dark photon models), the boundary conditions take the form,
@@
One can also derive a single second-order differential equation for the mode function $\mathcal{A}_{L}$, in full analogy with the transverse-polarizations case; cf. Eq. (III.1). However, in this case, it takes a rather cumbersome form, which we give in Appendix A [see Eq. (A) there] not to distract the reader from the main flow of the paper. One can use this equation for numerical calculations as well.
@@
From Eq. (74) we can find a reasonable definition of the momentum of the mode that crosses the horizon. In fact, when the second term in brackets becomes comparable to unity, the behavior of the mode starts deviating from the vacuum solution. Therefore, we define
@@
where again the upper monotonic envelope is used to ensure that once the mode crosses the horizon it never crosses it back during inflation. Further, for any given mode $k$, we can find a moment of time $t_{h,L}$ when it crosses the horizon,
@@ IV.2 Gradient-expansion formalism
Now, using either the definitions of functions $\mathcal{Q}^{(2p)}$, $\mathcal{F}^{(2p)}$, $\mathcal{K}^{(2p+1)}$ in coordinate space, Eqs. (87)–(89), or the spectral representations above, we use Eqs. (85)–(86) or, equivalently, Eqs. (66)–(67) in order to derive the equations of motion for the bilinear functions,
@@
The system of equations of the GEF is also infinite, as in the transverse case. Indeed, equations for $\mathcal{Q}^{(2p)}$ and $\mathcal{F}^{(2p)}$ contain $\mathcal{K}^{(2p+1)}$ while the equation for the latter contains $\mathcal{Q}^{(2p+2)}$ and $\mathcal{F}^{(2p+2)}$ etc. It can be truncated in two possible ways,
@@
Analogously to the case of transverse modes, the justification follows from the structure of the spectra of quadratic gauge-field quantities with a large number of spatial derivatives; see the discussion around Eq. (60).
@@ V Full system of equations
The quantities constructed from the magnetic field are the simplest ones since they contain only the transverse components,
@@
$=\frac{m^{4}a^{2}}{I_{1}}\langle A_{0}(-\triangle)^{-1}A_{0}\rangle=\frac{m^{2}}{I_{1}}\mathcal{Q}^{(-2)}\,$, (104)
and in the last line we used Eq. (15) for the divergence of the electric field.
Further, we express the terms quadratic in the vector potential. The square of the 0 component simply reads
@@
The square of the three-vector decomposes similarly to the electric field contribution,
@@ VI.1 Model specification
Initial conditions for the inflaton field are set at roughly 61 $e$-folds from the end of inflation (without gauge fields) by setting $\phi(0)=\phi_{0}=15.5\,M_{\rm P}$; the corresponding value for the inflaton derivative is taken from the slow-roll attractor solution $\dot{\phi}(0)\approx-V^{\prime}M_{\rm P}/\sqrt{3V}\big|_{\phi_{0}}=-\sqrt{2/3}m_{\phi}M_{\rm P}\approx-4.9\times 10^{-6}\,M_{\rm P}^{2}$.
@@
This allows us to explicitly omit all terms containing the $m^{2}/I_{1}$ factor in the Friedmann equations (V)–(V), Klein–Gordon equation (V) as well as in the GEF systems of equations (III.2)–(51) and (93)–(IV.2). Note that this does not correspond to switching to a massless vector field and discarding its longitudinal modes. We will see that if the mass is small but time dependent (in fact, inflaton dependent), there may be efficient vector-field production, especially of its longitudinal modes. In other words, although we are setting $\mu\ll 1$, we keep $M=\dot{m}/(mH)$ finite.
In the low-mass approximation above, the transverse GEF system (III.2)–(51) takes the form that is exactly the same as in the case of a massless gauge field. It was extensively studied in the literature previously, see Refs. [57, 58, 60, 65, 63]. Note that effects of time-dependent mass do not enter the set of equations for transverse modes; they are only sensitive to the kinetic and axial couplings. On the other hand, the longitudinal modes of the vector field are totally insensitive to the axial coupling as it never enters the corresponding equations of motion. Moreover, the kinetic coupling only enters the boundary terms while also not appearing in the GEF system (93)–(IV.2) explicitly. Since we are primarily interested in studying the production of longitudinal modes of comparable magnitude to the transverse ones, in our numerical analysis, we keep only the mass couplings to the inflaton and in some cases also the kinetic coupling. Thus, we always set $I_{2}(\phi)=0$ in the remaining part of the paper.
For the kinetic coupling, we assume the functional dependence proposed by Ratra [68] in the context of inflationary magnetogenesis,
@@
where $\beta$ is the dimensionless coupling constant of the model. In principle, in order to avoid the strong coupling problem, one should take only the values $\beta>0$. However, in our numerical analysis below, we will consider also some negative values just to see the qualitative differences in the picture of the vector-field production. [^3] In a realistic setup that avoids the strong coupling problem, one can accommodate the coupling function with negative $\beta$ at finite time intervals if they are followed by the intervals with $\beta>0$ (e.g., in a case of a piecewise coupling function); see, e.g., Refs. [69, 70, 71, 72, 73] for the case of inflationary magnetogenesis.
@@
The exponential-type functions as in Eqs. (113)–(114) can naturally arise in models where the vector field is nonminimally coupled to gravity. Following the seminal work [54], one can consider interactions of the form $R^{p}A_{\mu}A^{\mu}$ (as well as analogous couplings $R^{q}F_{\mu\nu}F^{\mu\nu}$ to the kinetic term). When such terms are embedded into $f(R)$ gravity or Higgs inflation and transformed to the Einstein frame, the scalaron or Higgs field plays the role of the inflaton and typically couples to the vector field through exponential functions. This mechanism has been explored, for example, in Refs. [74, 75, 76] for massless gauge fields, and its generalization to the massive case with a coupling $\sim R^{p}A_{\mu}A^{\mu}$ is straightforward.
Another, and arguably more natural, form of the mass coupling arises in models where the vector mass is generated through symmetry breaking, as in the Higgs mechanism; see, e.g., Refs. [51, 77, 78]. In this case, the mass typically scales as $m(\phi)\propto\phi$. For such a coupling one finds $M=\frac{\dot{\phi}}{H\phi}=\frac{\sqrt{2\epsilon}}{\phi/M_{\rm P}}\ll 1$, since $\epsilon=-\dot{H}/H^{2}\ll 1$ during slow-roll inflation and $\phi/M_{\rm P}\gtrsim 1$ in the chaotic inflation model (111) used in our numerical analysis. Therefore, gauge-field production is expected to be very similar to the case of a constant mass, i.e., $M=0$, which has already been studied in the literature (see, e.g., Ref. [32]). Although our method is fully capable of reproducing the gauge-field evolution in this scenario, no significant backreaction from the longitudinal modes is anticipated, and the result can be obtained straightforwardly using conventional mode-by-mode analyses in momentum space. For this reason, we do not include this type of coupling in our numerical study.
@@
Figure 1: Upper panel: the dependence of the energy densities $\rho_{\mathcal{Q}}$ (green dashed-dotted line) and $\rho_{\mathcal{F}}$ (purple dotted line) on the number of $e$-folds counted from the end of inflation in the model of mass coupling (114) of the vector field to inflaton with $\beta=12$, obtained by using the GEF truncated at the order $2p_{\rm max}+1=67$. Lower panel: evolution of the relative error of the GEF solution with respect to the mode-by-mode solution in momentum space.